If the area of a rectangle is 36mē. The length is 4mē. What is the perimeter?
Found 2 solutions by edjones, 2897696:
Answer by edjones(8007)

4w=36
w=9
P=4*2+9*2=26m
.
Ed

Answer by 2897696(96)

A rectangle has 2 lengths and 2 width
Starting from the area, we can find the missing dimensions, the formula in find area of rectangle is length* width.
plug numbers you have to formula.
36m^2=4m*x ==> let x be how long width is
simplify: 36=4x divide four on both sides
length of width is 9m
now you have the dimensions of the rectangle: width=9m length=4m
from here find the perimeter by adding the sides together.
there are 2 lengths and 2 with in a rectangle so the formula for perimeter:
P= w+w+l+l or simply P=2w+2L
plug into formula (either one)
P=9+9+4+4 or P=2(9)+2(4)
P=26m =======P= 26m either formula should get you the same answer.


Answer: perimeter of rectangle is 26m the missing side was 9m